F Question 4: Choosing the product mix when capacity is in short supply

You have three product lines: Basic, Premium, and Supreme. You have limited capacity of 300 machine hours. You face excess demand for all three products (assume unlimited demand for simplicity).

Basic Premium Supreme
Price per unit $1,800 $4,200 $6,000
Unit variable cost $600 $1,200 $1,800
Machine hours per unit 1 2 4


Compute the CM per unit of capacity (i.e, per machine-hour) for each product:
   Basic = $  per hour
   Premium = $  per hour
   Supreme = $  per hour

Which product(s) should you make?

Make all three products to take advantage of the high demand

Supreme only    

Premium only

Basic only


How many units of each product should you make? (enter 0 for products that you do not want to make)
   Basic =  units
   Premium =  units
   Supreme =  units
Hint if you get stuck: You have 300 machine hours.
Answer #1

1)contribution margin

contribution margin=p -v

p=price per product

v =variable cost per product

Basic=$1800-$600

=$1200 per hour

Premium==$4200-$1200

=$3000

machine hour=$3000/2

=$1500 per hour

Supreme=$6000-$1800

=$4200

machine hour=4200/4

=$1050 per hour

2)

premium only because it offers the higher level of contribution margin which will increase the profitability.

3)

Basic =0 unit

premium =

2 machine hour to create one unit

total machine hour= 300 hour

total quantity= 300/2

=150 Quantity

supreme =0 units